# High-throughput Parasitic-independent Probe Thermal Resistance Calibration for Robust Thermal Mapping with Scanning Thermal Microscopy

### v3

@@ Abstract
Nanostructured materials, critical for thermal management in semiconductor devices, exhibit a strong size dependence in thermal transport. Specifically, studying the variation of thermal resistance across grain boundaries is of critical importance in designing effective thermal interface materials for heterogeneous integration. Frequency-domain Thermoreflectance (FDTR)-based techniques can provide thermal resistance mapping in the micrometer (m) length scale. Scanning Thermal Microscopy (SThM) has the potential for quantification of local thermal transport with orders of magnitude higher spatial resolution (< 100 nm). However, challenges in the accurate quantification of the raw signal to thermal conductivity and surface sensitivity limit its widespread adoption as a characterization standard for understanding nanoscale heat transport and defect-mediated thermal properties modulation in nanostructured films. Here, we introduce a circuit-based probe thermal resistance () calibration technique that is independent of any parasitic heat transport pathways; thus, the calibration allows accurate determination of the heat dissipated from the SThM probe and the resulting tip temperature increase, thereby enabling extraction of the local thermal resistance of the film. In this study, SThM achieved sub-100 nm spatial resolution in mapping thermal resistance across a 15 nm-thick Al film deposited via e-beam evaporation on SiO2 substrate. Finally, the thermal resistance mapping is converted to thermal conductivity using the finite element modeling-based calibration technique, where the average of the pixel-level values yields the effective thermal conductivity of 2.87 0.18 W/mK, in good agreement with published theoretical frameworks explaining heat transport modes in ultra-thin Al.
@@ III.4 Modeling contact thermal resistance ()
As discussed in Sec. II, is an experimentally measured quantity that comprises the thermal properties of the sample. In addition, also includes thermal resistances arising from radiation (), air conduction (), and the water meniscus formed at the tip–sample contact (). The contribution of radiative heat transfer is often considered negligible, particularly under ambient SThM conditions with a cantilever–sample separation larger than 1 m li2025quantitative. The probe considered in our measurement has a tip height of 10 measured from the cantilever. In a humid environment, as shown in Fig. 2(b), a water meniscus forms at the tip–sample contact and serves as an additional pathway for heat transfer. A detailed study of water-meniscus formation and a theoretical model for heat transport through the meniscus were presented in Refs. gomes1999dc and luo1997sensor, respectively. Furthermore, Refs. zhang2020review and assy2015temperature reported that the contribution of the meniscus to the total tip–sample contact resistance () lies between 1 and 6. Therefore, the water-meniscus effect is treated as a 6 uncertainty in the values. Finally, under the assumption that , we express the measured quantity as the sum of three thermal resistances, as shown in Eq. (13).
Here, denotes the SThM tip thermal resistance when the tip is in contact; , the tip–sample interfacial thermal resistance; and , the sample spreading resistance.
@@
The impact of the interfacial thermal resistance () and its physically reasonable range are determined using FEM modeling, as discussed in Sec. IV. Furthermore, heat-spreading analysis is crucial for determining the thermal conductivity () of the sample. Yovanovich et al.yovanovich2002analytical and Muzychka et al.muzychka1999thermal derived analytical solutions for the thermal spreading resistance (), as shown in Eq. (14), for a film on a substrate as a function of the film thickness () and the thermal conductivities of the film () and substrate ().
where is defined as , is the source radius, and is the first-order Bessel function of the first kind. Furthermore, Muzychka and Yovanovich muzychka2004thermal derived a more compact expression for the spreading resistance of a circular isoflux heat source on a finite-thickness isotropic slab. Instead of evaluating the full integral in Eq. (14), the spreading resistance can be expressed in a closed correlation form as shown in Eq. (15).
where is a dimensionless correction function that accounts for finite substrate thickness and conductivity mismatch between the film and the substrate. This compact formulation reduces to the classical Mikic–Yovanovichyovanovich2005four result in the limit of a semi-infinite substrate ().
@@ IV.1 Finite-element multiphysics modeling of SThM probe
The effect of the on local sample heating was examined for two materials with strongly contrasting intrinsic thermal conductivities: Cu (300 W/mK) and SiO2 (1.4 W/mK). As shown in Fig. 4(b–d) for SiO2 and Fig. 4(f–h) for Cu, identical values of result in qualitatively similar surface temperature distributions beneath the tip, despite the large difference in bulk thermal conductivity. Because the COMSOL model considers smooth surfaces for all domains, the resulting temperature profiles for Cu and SiO2 are expected to be identical, which is in agreement with previous studies showing that is independent of the underlying material in the absence of surface roughness effects li2025quantitative; pernot2021frequency; bahrami2004thermal. At lower interfacial thermal resistance ( K/W), efficient heat transfer from the probe leads to a broader thermal diffusion profile in both materials. In contrast, increasing K/W strongly suppresses heat flow across the interface, resulting in negligible local surface heating for both Cu and SiO2. Importantly, the experimentally measured total thermal resistance lies in the range K/W and is accompanied by clear local surface heating beneath the probe. Therefore, must be K/W and cannot exceed the experimentally inferred thermal resistance.
Furthermore, the difference between the SThM tip temperature and the corresponding local sample temperature variation is plotted in Fig. 5(b). The change in tip and sample temperature () is shown for various values. The results shown in Fig. 5(a,b) demonstrate that plays a dominant role in the quantitative determination of the sample’s local thermal conductivity. While is treated parametrically in the FEM simulations to examine its influence on the measured temperature response, this approach is adopted to establish a physically constrained relationship between the experimentally accessible temperature contrast and . By comparing the experimentally measured (Fig. 3) with the trends predicted by the FEM model shown in Fig. 5 (b), a plausible and bounded range of is identified for the given probe–sample contact conditions. Although the thermal conductivity extraction model used in this study does not explicitly depend on the absolute value of , the FEM model serves as a sanity check and establishes a physically reasonable range for .
@@ IV.2 Experimental determination of model parameters
The thermal conductivity of Al is determined using the procedure described in Refs. huang2024violation and spiece2021quantifying, which reduces noise and measurement uncertainty. To extract the thermal resistance associated with Al and SiO2, a pixel-to-pixel correlation between the height and thermal signal maps is performed. This correlation is used to construct a 2D histogram (see Fig. S2 from the Supplementary Material), which enables a separation of the thermal responses from Al and SiO2 surfaces. Traditional analysis methods often rely on arbitrarily selected line-cuts or cropped regions of interest, introducing potential user bias in the selection process. In contrast, the present approach systematically processes all available data points from the thermal signals measurement, ensuring reproducibility and minimizing subjective interpretation. This comprehensive analysis leverages the full spatial information content, thereby improving statistical confidence in the extracted thermal properties. Subsequently, the acquired data are subjected to statistical processing to extract thermal properties. However, this analysis is susceptible to errors arising from experimental artifacts—such as surface contamination, tip-sample contact variations, or instrumental noise—which can distort the thermal signal and compromise measurement accuracy. Therefore, to quantitatively extract representative thermal resistance values while accounting for spatial heterogeneity, a 2D histogram of thermal resistance on the y-axis and film thickness on the x-axis is constructed (Fig. S2 of Supplementary Material). This histogram is then projected onto the thermal resistance axis to obtain statistically representative values. This projection combines measurements across all thickness values, yielding a one-dimensional distribution from which statistically robust resistance values can be determined. The resulting 1D distribution is modeled using a double-Gaussian function, from which the mean peak positions corresponding to the thermal resistances of Al and SiO2 are extracted. The statistical uncertainties in these mean values are obtained from the covariance matrix of the nonlinear least-squares fitting procedure. From this analysis, the mean thermal resistances were found to be 2.60 0.03 MK/W and 5.41 0.04 MK/W for Al and SiO2, respectively.
In the following analysis, the thermal resistance distributions presented in Figure 6 (e,f) are utilized to quantitatively determine the thermal conductivity of 15 nm Al thin film. Following the theoretical framework established with Eq. (13) in Sec. III, we model the measured thermal resistance as a series combination of the tip thermal resistance (), sample spreading resistance (), and the film-substrate interfacial thermal resistance () as previously described in Ref. [huang2024violation]. This formulation enables direct extraction of the Al thermal conductivity from the experimental data. Therefore, the measured thermal resistance () consists of three components: , , and . Among these contributions, as calculated in Sec. III(C), K/W, which is at least two orders of magnitude smaller than both and . Furthermore, because the thermal conductivity is evaluated relative to the SiO2 reference surface, the contribution of can be neglected.
@@
Furthermore, since the experimentally measured thermal resistance () values are K/W, it can be concluded that is also K/W, as discussed in Sec. IV(A). In addition, the tip–sample contact resistance () is mainly governed by surface roughness, the applied tip–sample normal force, and the resulting contact area pernot2021frequency; bahrami2004thermal; puyoo2010thermal. Therefore, the thermal contact between the SThM probe and the sample may slightly differ for Al and SiO2 surfaces; however, it is very challenging to determine the exact for each case. Therefore, for simplification, we assumed the same for both Al and SiO2 interfaces. This assumption introduces negligible uncertainty in the measurement because both surfaces exhibit comparable roughness (see Fig. S6 in the Supplementary Material). Additionally, the tip-sample normal contact force was maintained with an AFM setpoint value of 0.6 V, which corresponds to < 10 nN contact force for this type of probes. This assumption allows us to isolate the spreading resistance component, which leads to the measured thermal resistance () on the Al film surface. Finally, the thermal conductivity is calculated using Eq. (15), under the assumption of a semi-infinite substrate. The analysis of the statistical distribution of thermal resistance values yields W/mK for the 15 nm Al film. Detailed calculation of is provided in S1 of Supplementary Material. We assumed a literature value for the thermal conductivity of 1.38 W/mK for the SiO2 substrate grove1967physics. The reported uncertainty accounts for both statistical variations and thermal signal errors arising from additional heat transfer through the water meniscus formed around the SThM tip assy2015temperature.
V Discussion
| Metal | (W/m·K) | (W/m·K) | (W/m·K) | Refs. |
@@ V Discussion
| Al | 6 | 246 | 237–252 | [jain2016thermal, kittel2018introduction] |
@@
The thermal conductivity of Al consists of both phonon () and electron () contributions, with electrons typically dominating heat transport in bulk. Jain et al. performed detailed first-principles calculations to examine both phonon-phonon and electron-phonon scattering processes in bulk Al, quantitatively determining the individual contributions of and to the overall thermal conductivity jain2016thermal. This study further quantified the mean free path () distributions of both phonon and electron heat carriers by analyzing their respective thermal conductivity accumulation functions at 300 K (see Fig. S7 of Supplementary Material). Analysis of the thermal conductivity accumulation function demonstrates that the electron contribution decreases substantially as decreases, indicating that electrons with short contribute less to heat transport than those with longer . This suggests that long-range electron transport dominates thermal conductivity in bulk Al, whereas phonon transport shows the inverse behavior, being dominated by the short scattering processes.
@@
The accumulation function data at 300 K indicate that phonon thermal conductivity in Al is dominated by carriers with mean free paths ranging from 1 nm to 22 nm (comprising of ), while electron thermal conductivity requires longer mean free paths of 10 nm to 25 nm for similar accumulation ( of ). The onset of contributions also differs markedly: phonon transport begins at , whereas electron transport only becomes substantial at . These characteristic length scales directly impact thermal transport in confined geometries. In our 15 nm Al film, a sub-stoichiometric Al2O3-x layer (1–2 nm thick) forms naturally on the Al surface due to exposure to ambient air yang2018liquid; gorobez2021growth, reducing the effective conductive metallic thickness to approximately 13–14 nm. This confinement preferentially suppresses electron transport, as the film dimension is smaller than the typical electron mean free paths, resulting in diffuse boundary scattering that significantly reduces . Conversely, phonons with mean free paths of 1–13 nm can propagate through the film with less boundary interference. This size-dependent scattering leads to a crossover in the dominant heat carrier: while bulk Al thermal conductivity is electron-dominated, phonon transport becomes the primary mechanism in this nanoscale thin film. Therefore, the thermal conductivity obtained in this study ( W/mK) falls within a reasonable and expected range for nanoscale Al films. This value, approximately 80 times lower than bulk Al, is consistent with the dominance of phonon transport and the suppression of electron contributions due to dimensional confinement.
@@
The thermal signal and topography appear elevated along the Al/SiO2 boundary (see Fig. 6(a, c)). The elevated topography introduces a minor artifact in the thermal signal due to AFM feedback sensitivity effects at the step height. Consequently, the higher thermal resistance extracted at the edge arises from the same artifact, as it is directly derived from the thermal signal. Nevertheless, Fig. 6(e) demonstrates the significant potential of this approach for probing thermal transport variations across dislocation defects and grain boundaries with sub-100 nm spatial resolution. Furthemore, validation using established approaches, such as those described in Refs. tiwari2013anisotropic; isotta2023microscale for assessing thermal property variations across grain boundaries with different densities, would be highly beneficial, particularly given the high spatial resolution offered by SThM.
VI Conclusion

### v4

@@ Abstract
Nanostructured materials, critical for thermal management in semiconductor devices, exhibit a strong size dependence in thermal transport. Studying thermal resistance variation across grain boundaries is critical for designing effective thermal interface materials. Frequency-domain Thermoreflectance (FDTR)-based techniques can provide thermal resistance mapping at the micrometer (m) scale. Scanning Thermal Microscopy (SThM) enables quantification of local thermal transport with significantly higher spatial resolution ( nm). However, challenges in quantifying the raw signal to thermal conductivity and surface sensitivity limit its widespread adoption for understanding nanoscale heat transport and defect-mediated thermal properties in nanostructured films. Here, we introduce a circuit-based probe thermal resistance () calibration technique independent of parasitic heat pathways, enabling accurate determination of probe heat dissipation and tip temperature rise, thereby allowing extraction of local thermal resistance. SThM achieved sub-100 nm spatial resolution in mapping thermal resistance across a 15 nm-thick Al film deposited via e-beam evaporation on SiO2 substrate. The thermal resistance maps are converted to thermal conductivity using robust analytical and finite element models that account for tip-sample geometry, lateral heat spreading, and buried interface effects. Gaussian distribution fitting of pixel-level thermal resistance values yields W/(mK) for the ultra-thin Al film (13–15 nm), representing a 5.3-fold reduction from bulk aluminum (237 W/(mK)). These results agree with published experimental data and theoretical frameworks explaining thickness-dependent heat transport in ultra-thin metallic films.
@@ III.4 Modeling contact thermal resistance ()
As discussed in Sec. II, is an experimentally measured quantity that comprises the thermal properties of the sample. In addition to solid-solid thermal resistance (), also includes thermal resistances arising from radiation (), air conduction () and the water meniscus formed at the tip–sample contact (). The contribution of radiative heat transfer is often considered negligible, particularly under ambient SThM conditions with a cantilever–sample separation larger than 1 m li2025quantitative. The probe considered in our measurement has a tip height of 10 measured from the cantilever. In a humid environment, as shown in Fig. 2(b), a water meniscus forms at the tip–sample contact and serves as an additional pathway for heat transfer. A detailed study of water-meniscus formation and a theoretical model for heat transport through the meniscus were presented in Refs. gomes1999dc and luo1997sensor, respectively. Furthermore, Refs. zhang2020review and assy2015temperature reported that the contribution of the meniscus to the total tip–sample contact resistance () lies between 1 and 6. Therefore, the water-meniscus effect is treated as a 6 uncertainty in the values. Therefore, under the assumption that , we express the measured quantity as the sum of three thermal resistances, as shown in Eq. (13).
where:
: Tip-sample interface thermal resistance
: Spreading resistance (Yovanovich Eq. (14))yovanovich2002analytical; muzychka1999thermal
: Interface boundary resistance
: Geometric dilution factor
: Lateral spreading length
@@
The impact of the interfacial thermal resistance () and its physically reasonable range are determined using FEM modeling, as discussed in Sec. IV. Furthermore, heat-spreading analysis is crucial for determining the thermal conductivity () of the sample. Yovanovich et al.yovanovich2002analytical; yovanovich2005four and Muzychka et al.muzychka1999thermal derived analytical solutions for the thermal spreading resistance (), as shown in Eq. (14), for a film on a substrate as a function of the film thickness () and the thermal conductivities of the film () and substrate ().
where is defined as , is the source radius, and is the first-order Bessel function of the first kind.
The buried interface thermal boundary conductance is a critical parameter in this analysis. However, as heat transfers through a thin film to the substrate, it spreads laterally before reaching the buried interface. The characteristic lateral spreading length () determines the effective area over which heat crosses the buried interface. For heat flow from a localized source of radius at the top of a thin film of thickness and thermal conductivity , deposited on a substrate with thermal conductivity , this spreading length is given by Eq. (15)muzychka2004thermal; yovanovich2003thermal; kennedy1960spreading.
This lateral spreading reduces the effective contribution of the buried interface resistance through a geometric dilution effect. The dilution factor represents the fraction of interface resistance “seen” by the heat flux and is given by Eq. (16).
Physically, heat spreading reduces the effective contribution of interface resistance by the ratio of the original contact area to the expanded area at the interface. As the spreading length increases relative to the contact radius, more heat bypasses the localized interface resistance, thereby reducing its impact on the total thermal resistance.
A critical insight is that both and depend on , which is the unknown quantity we seek to determine. This creates a self-consistency requirement: the spreading resistance and the effective buried interface resistance are both functions of the unknown film thermal conductivity. Consequently, the thermal resistance balance (Eq. 17)
must be solved iteratively to extract . This iterative solution converges to yield the thermal conductivity of the thin film, with the complete algorithm detailed in the section S2. (Supplementary Material).
@@ IV.1 Finite-element multiphysics modeling of SThM probe
The effect of the on local sample heating was examined for two materials with strongly contrasting intrinsic thermal conductivities: Cu (300 W/mK) and SiO2 (1.38 W/mK). As shown in Fig. 4(b–d) for SiO2 and Fig. 4(f–h) for Cu, identical values of result in qualitatively similar surface temperature distributions beneath the tip, despite the large difference in bulk thermal conductivity. Because the COMSOL model considers smooth surfaces for all domains, the resulting temperature profiles for Cu and SiO2 are expected to be identical, which is in agreement with previous studies showing that is independent of the underlying material in the absence of surface roughness effects li2025quantitative; pernot2021frequency; bahrami2004thermal. At lower interfacial thermal resistance ( K/W), efficient heat transfer from the probe leads to a broader thermal diffusion profile in both materials. In contrast, increasing K/W strongly suppresses heat flow across the interface, resulting in negligible local surface heating for both Cu and SiO2. Importantly, the experimentally measured total thermal resistance lies in the range K/W and is accompanied by clear local surface heating beneath the probe. Therefore, must be K/W and cannot exceed the experimentally inferred thermal resistance.
Furthermore, the difference between the SThM tip temperature and the corresponding local sample temperature variation is plotted in Fig. 5(b). The change in tip and sample temperature () is shown for various values. The results shown in Fig. 5(a,b) demonstrate that plays a dominant role in the quantitative determination of the sample’s local thermal conductivity. While is treated parametrically in the FEM simulations to examine its influence on the measured temperature response, this approach is adopted to establish a physically constrained relationship between the experimentally accessible temperature contrast and . By comparing the experimentally measured (Fig. 3) with the trends predicted by the FEM model shown in Fig. 5 (b), a plausible and bounded range of is identified for the given probe–sample contact conditions.
@@ IV.2 Experimental determination of model parameters
To extract the thermal resistance associated with Al and SiO2, a pixel-to-pixel correlation between the height and thermal signal maps is performed. This correlation is used to construct a 2D histogram (see Fig. S2 from the Supplementary Material), which enables a separation of the thermal responses from Al and SiO2 surfaces. Traditional analysis methods often rely on arbitrarily selected line-cuts or cropped regions of interest, introducing potential user bias in the selection process. In contrast, the present approach systematically processes all available data points from the thermal signals measurement, ensuring reproducibility and minimizing subjective interpretation. This comprehensive analysis leverages the full spatial information content, thereby improving statistical confidence in the extracted thermal properties. Subsequently, the acquired data are subjected to statistical processing to extract thermal properties. However, this analysis is susceptible to errors arising from experimental artifacts—such as surface contamination, tip-sample contact variations, or instrumental noise—which can distort the thermal signal and compromise measurement accuracy. Therefore, to quantitatively extract representative thermal resistance values while accounting for spatial heterogeneity, a 2D histogram of thermal resistance on the y-axis and film thickness on the x-axis is constructed (Fig. S2 of Supplementary Material). This histogram is then projected onto the thermal resistance axis to obtain statistically representative values. This projection combines measurements across all thickness values, yielding a one-dimensional distribution from which statistically robust resistance values can be determined. The resulting 1D distribution is modeled using a double-Gaussian function, from which the mean peak positions corresponding to the thermal resistances of Al and SiO2 are extracted. The statistical uncertainties in these mean values are obtained from the covariance matrix of the nonlinear least-squares fitting procedure. From this analysis, the mean thermal resistances were found to be 2.60 0.03 MK/W and 5.41 0.04 MK/W for Al and SiO2, respectively.
The measured thermal resistances MK/W and MK/W provide the experimental foundation for quantitative determination of the thermal conductivity of the Al thin film. Following the theoretical framework established in Sec. III, we model the measured contact-to-noncontact thermal resistance difference as a series combination of three contributions: the tip-sample interfacial thermal resistance (), the sample spreading resistance (), and the Al/SiO2 buried interface thermal boundary resistance (), as given by Eq. (13). This formulation enables direct extraction of the Al thermal conductivity from the experimental data through an iterative algorithm (Eq. 17) that accounts for the interdependence of and on the unknown film thermal conductivity.
IV.2.1 Tip Interface Resistance Extraction
@@
As shown in Fig. 6(e), the experimentally measured values are K/W. Therefore, it can be concluded that is also K/W. In addition, the tip–sample contact resistance () is mainly governed by surface roughness, the applied tip–sample normal force, and the resulting contact area pernot2021frequency; bahrami2004thermal; puyoo2010thermal. Therefore, the thermal contact between the SThM probe and the sample may slightly differ for Al and SiO2 surfaces; however, it is very challenging to determine the exact for each case. Therefore, for simplification, we assumed the same for both Al and SiO2 interfaces. This assumption introduces negligible uncertainty in the measurement because both surfaces exhibit comparable roughness (see Fig. S6 in the Supplementary Material). Additionally, the tip-sample normal contact force was maintained with an AFM setpoint value of 0.6 V, which corresponds to < 10 nN contact force for this type of probes. Therefore, the tip-sample interface resistance is extracted from a reference measurement on bare SiO2/Si (no thin film, ).
where is calculated using Eq. (14) with known (1.38 W/mK) and (150 W/mK). The parameters used in this analysis are provided in Table S1 (supplementary material).
IV.2.2 Iterative Algorithm for Extraction
Literature values for metal/dielectric interfacesstoner1993kapitza; lombard2014influence; schmidt2008pulse; cheaito2015thermal span a considerable range. For the Al/SiO2 interface, we adopt a central value of MW/(m2K) with a realistic range of –250 MW/(m2K) to account for this uncertainty hopkins2010criteria; kwon2021thermal; zhu2010ultrafast.
Applying the iterative algorithm to the experimental measurements yields a thermal conductivity of W/(mK) at the central values of MW/(m2K) and nm. See Section S3 (supplementary material) for the detailed calculation. In our 15 nm Al film, a sub-stoichiometric Al2O3-x layer (1–2 nm thick) forms naturally on the Al surface due to exposure to ambient airyang2018liquid; gorobez2021growth, reducing the effective conductive metallic thickness to approximately 13–14 nm. A comprehensive uncertainty analysis accounting for variations in both interface conductance (–250 MW/(m2K)) and film thickness (–15 nm) produces asymmetric error bounds of W/(mK), yielding a final result of W/(mK). This represents a 5.3-fold reduction from bulk aluminum (237 W/(mK)), consistent with Fuchs-Sondheimer theoryfuchs1938conductivity; Sondheimer1952MeanFreePath for films of thickness comparable to the electron mean free path (). The thermal resistance budget (Table 2) reveals that the tip-sample interface dominates the total measured resistance (70.8%), with spreading resistance contributing 26.2% and the diluted buried interface resistance only 3.1%. The geometric dilution factor of indicates that lateral heat spreading over a characteristic length of nm substantially reduces the effective contribution of the Al/SiO2 interface resistance. A complete matrix of extracted thermal conductivity values for all combinations of and thickness is provided in the Table. S1 of Supplementary Material.
| Component | Value (MK/W) | Percentage | Description |
| (tip-Al) | 1.84 | 70.8% | Tip-sample interface |
| (Al/SiO2) | 0.68 | 26.2% | Spreading resistance |
| (Al/SiO2) | 0.08 | 3.1% | Diluted |
| Total | 2.60 | 100% | = |
| Metal | (W/mK) | (W/mK) | (W/mK) |
| Al | 6 | 246 | 237–252 |
@@ IV.2.2 Iterative Algorithm for Extraction
Data adapted from Refs. jain2016thermal; kittel2018introduction.
Discussion
@@
The thermal conductivity of Al consists of both phonon () and electron () contributions, with electrons typically dominating heat transport in bulk. Jain et al. performed detailed first-principles calculations to examine both phonon-phonon and electron-phonon scattering processes in bulk Al, quantitatively determining the individual contributions of and to the overall thermal conductivityjain2016thermal. Furthermore, they quantified the mean free path () distributions of both phonon and electron heat carriers by analyzing their respective thermal conductivity accumulation functions at 300 K (see Fig. S7 of Supplementary Material). Analysis of the thermal conductivity accumulation function demonstrates that the electron contribution decreases substantially as decreases, indicating that electrons with short contribute less to heat transport than those with longer .
@@
The accumulation function data at 300 K indicate that phonon thermal conductivity in Al is dominated by carriers with mean free paths ranging from 1 nm to 22 nm (comprising 90% of ), while electron thermal conductivity requires longer mean free paths of 10 nm to 25 nm for similar accumulation (90% of ). The onset of contributions also differs markedly: phonon transport begins at nm, whereas electron transport only becomes substantial at nm. These characteristic length scales directly impact thermal transport in confined geometries. This confinement preferentially suppresses electron transport, as the film dimension is smaller than the typical electron mean free paths, resulting in diffuse boundary scattering that significantly reduces .
Fig. 7 compares the SThM-extracted thermal conductivity of the 14 nm Al film ( W/(mK)) with literature values for polycrystalline Al thin films of varying thickness measured by different techniques, including frequency-domain thermoreflectance (FDTR) Schmidt2010ThinMetalFDTR, photothermal mirage Wu1993PhotothermalThinFilms, electrical heating with temperature profile analysis lugo2016thermal, and electron diffraction thermometry Volkov1976AluminumThermal. Boltzmann transport equation (BTE) modeling results from Jin et al. Jin2008ElectronMeanFreePath are also included. Our result falls on the expected trend of decreasing thermal conductivity with decreasing film thickness, consistent with enhanced electron scattering at surfaces and grain boundaries.
The dashed curve in Fig. 7 is a trend line that captures the thickness-dependent behavior and remains consistent with the combined Fuchs-Sondheimer and Mayadas-Shatzkes electron scattering model fuchs1938conductivity; Sondheimer1952MeanFreePath; mayadas1970electrical. Using the literature value of nm for the electron mean free path in aluminum Gall2016, this theoretical framework predicts W/(mK) for a 14 nm film (Supplementary Section S4), which falls within our experimental uncertainty range. The spread in literature data reflects variations in film microstructure, deposition conditions, and measurement techniques across different studies. The agreement between our SThM measurement, the theoretical prediction, and existing literature data validates the iterative extraction methodology developed in this work.
@@
The thermal signal and topography appear elevated along the Al/SiO2 boundary (see Fig. 6(a, c)). The elevated topography introduces a minor artifact in the thermal signal due to AFM feedback sensitivity effects at the step height. Consequently, the higher thermal resistance extracted at the edge arises from the same artifact, as it is directly derived from the thermal signal. Nevertheless, Fig. 6(e) demonstrates the significant potential of this approach for probing thermal transport variations across dislocation defects and grain boundaries with sub-100 nm spatial resolution. Furthermore, validation using established approaches, such as those described in Refs. tiwari2013anisotropic; isotta2023microscale for assessing thermal property variations across grain boundaries with different densities, would be highly beneficial, particularly given the high spatial resolution offered by SThM.
V Conclusion